OpenStudy (anonymous):

Twenty years ago a small town in Texas had a population of 10,000. The population has increased 8% each year since then. What is the growth factor that would be used to solve this problem? a. The growth factor would be 8 b. The growth factor would be 0.08 c. The growth factor would be 1.08 d. The growth factor would be 0.92

OpenStudy (triciaal):

percent means per 100 8% = 8/100 = 0.08

OpenStudy (anonymous):

can someone please help me!

OpenStudy (anonymous):

thank you so much!

OpenStudy (triciaal):

to get the population it would be the original PLUS the increase to get the new population can use original * 1.08

OpenStudy (anonymous):

Find the missing value for the exponential function represented by the table below. X. -2 Y. 21 -1 14.7 0 10.29 1 2 5.0421 a. 9.213 c. 7.203 b. 8.756 d. -7.203

OpenStudy (anonymous):

can you explain it to me @triciaal

OpenStudy (triciaal):

multiple choice I would use C I don't like this question the factor should just be by how much it has increased 0.08 as stated but to find the new population use 1.08 original plus increase

OpenStudy (anonymous):

can you explain the one above pls

OpenStudy (triciaal):

do you have the formula?

OpenStudy (triciaal):

is it y = e^x?

OpenStudy (anonymous):

no all i have is what i sent you. i need to find the missing value for X

OpenStudy (triciaal):

can you use logs?

OpenStudy (anonymous):

no it doesnt have a formula

OpenStudy (triciaal):

@jim_thompson5910 @satellite73 can you please help

OpenStudy (triciaal):

@amistre64 , @acxbox22 can you please help

OpenStudy (triciaal):

@besqueda9 sorry I don't know. tried to get you some help

OpenStudy (anonymous):

thank you so much anyway. youve been helpful

OpenStudy (triciaal):

@bibby

jimthompson5910 (jim_thompson5910):

hint: notice how sqrt(21*10.29) = 14.7 the 21 and 10.29 are values in the y column. The 14.7 is between 21 and 10.29

OpenStudy (amistre64):

you will have to develop an exponential function using a system of equations y = a b^(x-h) + k

OpenStudy (amistre64):

not 'have to' but that would be my approach

OpenStudy (triciaal):

@jim_thompson5910 , @amistre64 thank you